OpenStudy (anonymous):

two lines are perpendicular. the equation of the first line is y=4x+1. which of the following CANNOT be the equation of the second line.

OpenStudy (anonymous):

\[y=-\frac{ 1 }{ 4 }x-2\]

OpenStudy (anonymous):

\[x-4y=-2\]

OpenStudy (anonymous):

\[x+4y=-2\]

OpenStudy (anonymous):

\[8y=-2x-5\]

OpenStudy (anonymous):

@Ethanxiaorox

OpenStudy (anonymous):

i put them all into that calculator thing and they all intersect ? doesnt that make them all perpendicular

OpenStudy (animeasgardian):

I did it too...that's weird

OpenStudy (anonymous):

One second, my moms calling me :P if i take too long u can tag TheSmartONe, Or Igreen

OpenStudy (anonymous):

is that what perpendicular means? just to like intersect?

OpenStudy (animeasgardian):

yeah

OpenStudy (anonymous):

wtf

OpenStudy (animeasgardian):

ikr...

OpenStudy (animeasgardian):

double check to make sure you didn't type any of the equations wrong

OpenStudy (anonymous):

Im back .-.

OpenStudy (anonymous):

so i think its the x-4y=-2 cause i put them all in at once and that is the only line that looks different than the other ones

OpenStudy (anonymous):

okay ahah

OpenStudy (anonymous):

Im back :D My mom gave me a pice of cake :P But it has strawberiies .-.

OpenStudy (anonymous):

how sweet of your mom

OpenStudy (anonymous):

So did you get an answer?

OpenStudy (anonymous):

im pretty sure its the one i said

OpenStudy (anonymous):

i dont know how to do this question tho now

OpenStudy (anonymous):

two perpendicular lines intersect at the origin. if the slope of the first line is \[\frac{ -1 }{ 2 }\], what is the equation of the second line?

OpenStudy (anonymous):

Wut? @TheSmartOne Help ples?

OpenStudy (anonymous):

Ill start working on it .-.

OpenStudy (anonymous):

oh yeah do you want me to put the choices?

OpenStudy (anonymous):

Yes, that would help

OpenStudy (anonymous):

a. \[y=-2x\]

OpenStudy (anonymous):

b. \[y=\frac{ 1 }{ 2 }x\]

OpenStudy (anonymous):

c. \[y=-\frac{ 1 }{ 2 }x\]

OpenStudy (anonymous):

d. \[y=2x\]

OpenStudy (anonymous):

I have no idea >.< Never leanred whatever frac is >.<

OpenStudy (anonymous):

ill chose a lol

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

@SolomonZelman @chosenmatt A little help please?

OpenStudy (anonymous):

D: